David needs R30835,42 to buy a second hand trailer. How long will it take him to save towards this amount if he deposits R25000,00 now into a savings account earning 10,5% interest per year compounded weekly
1
Expert's answer
2021-10-06T00:28:37-0400

In this we need to calculate number of weeks required to achieve that.

Deposit today=$25000

Future value=R30835.42

Interest rate=10.5%

"Weekly\space interest\space rate=\frac{10.5}{52}=0.2019230769\%\\

Future \space value= Deposit\space today (1+r)^n\\

30835.42=25000(1.0020192307)^n\\

1.2334168=(1.0020192307)^n\\

n=104\space weeks\\

n=2 \space years"

It will take 2 years to achieve that target.